# HumanGif: Single-View Human Diffusion with Generative Prior

### v1

@@ 3.2. Single-view conditioned Human Diffusion Model
SHERF (Hu et al., 2023) 9.75E-06 26.128 0.934 0.063 7.48E-06 27.435 0.946 0.048
Animate Anyone (Hu, 2024) 2.32E-05 21.022 0.929 0.064 1.35E-05 24.306 0.946 0.041 -
Champ (Zhu et al., 2024) 2.42E-05 21.326 0.930 0.064 1.34E-05 25.381 0.952 0.037 -
@@
SHERF (Hu et al., 2023) 3.77E-06 27.581 0.948 0.049 3.99E-06 27.471 0.944 0.048 12.01
Animate Anyone (Hu, 2024) 6.20E-06 23.499 0.902 0.056 6.84E-06 23.043 0.907 0.061 16.22
Champ (Zhu et al., 2024) 6.37E-06 23.715 0.869 0.054 6.96E-06 23.253 0.879 0.058 16.89
@@
Human NeRF. The proposed framework is supposed to synthesize consistent novel views and poses from a single human image. However, we observe that the learned human diffusion models encounter two challenges: a) difficulty in capturing fine-grained details (e.g., logos on T-shirts) even when such details are present in the reference image, and b) limited ability to ensure multi-view and multi-pose consistency. One underlying reason is that the reference image is first encoded into a latent representation and then processed by ReferenceNet to extract features, leading to information loss during this process. To tackle this challenge, drawing inspiration from previous research in Human NeRF (Peng et al., 2021c; Hu et al., 2023; Chen et al., 2021; Peng et al., 2021a; Xu et al., 2021; Noguchi et al., 2021; Weng et al., 2022b; Su et al., 2021; Jiang et al., 2022a, c; Wang et al., 2022; Kwon et al., 2021; Gao et al., 2022; Choi et al., 2022; Zhao et al., 2021; Huang et al., 2022), we introduce a Human NeRF module in human diffusion models to learn the fine-grained features spatially aligned with the reference and target pose space. As illustrated in Fig. 2, the input to the Human NeRF module is a single human image $\mathbf{I}^{\text{ref}}$ along with its corresponding camera parameters $\bm{P}^{\text{ref}}$ and SMPL pose $\bm{\theta}^{\text{ref}}$ and shape $\bm{\beta}^{\text{ref}}$. The module outputs the rendered human feature image in the target camera view $\bm{P}^{\text{tgt}}$, corresponding to the target SMPL pose $\bm{\theta}^{tgt}$ and shape $\bm{\beta}^{\text{tgt}}$. To render features in the target space, points $\bm{x}^{\text{tgt}}$ are sampled along the cast rays. These points $\bm{x}^{\text{tgt}}$ are then transformed into the canonical space $\bm{x}^{c}$ using inverse Linear Blend Skinning (LBS). Subsequently, hierarchical 3D-aware features,i.e., point-level features and pixel-aligned features, are queried from their respective feature extraction modules. The queried features are concatenated and passed to the NeRF decoder to predict the density $\sigma$ and feature $\bm{c}$. The final pixel features are rendered in the target space through volume rendering, integrating the density and feature values of the sampled 3D points along the rays in the target space. The details of extracting point-level features and pixel-aligned features are described as follows. 1) Point-level Features. We extract per-point features by projecting the SMPL vertices onto the 2D feature map of the input image, as illustrated in Fig. 2. Next, we apply inverse LBS to transform the posed vertex features into the canonical space. These transformed features are then voxelized into sparse 3D volume tensors and further processed using sparse 3D convolutions (Contributors, 2022). From the encoded sparse 3D volume tensors, we extract point-level features $\bm{f}_{\text{point}}(\bm{x}^{c})$ for each point $\bm{x}^{c}$. With the awareness of 3D Human structure, point-level features capture local texture details in the seen area and infer textural information for unseen areas through sparse convolution. 2) Pixel-aligned Features. Due to limited SMPL mesh and voxel resolution, point-level features suffer from significant information loss, especially in areas visible from the reference image. To compensate for the information loss problem, we additionally extract pixel-aligned features by projecting 3D deformed points $\bm{x}^{c}$ into the input view. As illustrated in Fig. 2, each deformed point $\bm{x}^{c}$ is transformed into the observation space as $\bm{x}^{\text{ref}}=\text{LBS}(\bm{x}^{c};\bm{\theta}^{\text{ref}},\bm{\beta}^{\text{ref}})$ using LBS. It is then projected onto the input view, allowing us to query the pixel-aligned features. Leveraging the complementary strengths of point-level and pixel-aligned features, our Human NeRF module effectively captures fine-grained feature details in regions visible in the reference view while inferring features for regions occluded in the reference view.
@@
Multi-view/Temporal Module. To capture multi-view and temporal consistency among a target image sequence, we further integrate a multi-view and temporal layer after the spatial-attention and cross-attention components within the Res-Trans block of Denosing UNet. The multi-view attention layer shares the same architecture as the temporal layer used in (Hu, 2024; Guo et al., 2023).
3.3. Training Detail
@@
(4) $\mathcal{L}=\mathcal{L}_{\text{diff-latent}}+\lambda_{1}\mathcal{L}_{\text{diff-img}}+\lambda_{2}\mathcal{L}_{\text{NeRF}},$
@@
To bridge the gap between latent space and image space (Müller et al., 2023), we formulate $\mathcal{L}_{\text{diff-img}}$ with approximation $\bm{\tilde{z}}_{0}(\bm{\epsilon},\Phi):=\bm{z}_{0}+\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}(\bm{\epsilon}-\bm{\epsilon}_{\Phi}(\bm{z}_{t},t,\bm{c}))$,
(5) $\mathcal{L}_{\text{diff-img}}=w_{t}\mathbb{E}_{\bm{z}_{t},\bm{\epsilon},t,\bm{c}}\Arrowvert\mathcal{D}(\bm{\tilde{z}}_{0}(\bm{\epsilon},\Phi))-\bm{I}^{\text{tgt}}\Arrowvert^{2},$
@@
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, $\bm{I}^{\text{tgt}}$ is the image at the target view with the target human pose, $\mathcal{D}$ denotes the decoder in a VAE (Kingma, 2013; Van Den Oord et al., 2017) model. Since the approximation holds reliably only for steps $t$ close to zero, we introduce a weight $w_{t}$ that progressively decays as the step value increases. In addition to MSE loss, we apply structural similarity index (SSIM) (Wang et al., 2004) and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) as additional image-level loss terms. For features from Human NeRF, we follow (Chan et al., 2022) to add a term $\mathcal{L}_{\text{NeRF}}$, which regularizes the first three channels by computing MSE, SSIM, and LPIPS with the target image $\bm{I}^{\text{tgt}}$ and a Mask Loss with the target human mask.
4. Experiments
Figure 3. Qualitative results of novel view synthesis (1st and 3rd row) and novel pose synthesis (2nd and 4th row) produced by SHERF, Animate Anyone, Champ and our HumanGif on RenderPeople and DNA-Rendering datasets.
4.1. Experimental Setup
@@
Datasets. We evaluate the performance of our HumanGif on two multi-view human modelling datasets, i.e., RenderPeople (ren, 2018) (Synthetic Humans) and DNA-Rendering (Cheng et al., 2023) (Real-World Humans). For RenderPeople, we randomly sample 450 subjects as the training set and 30 subjects for testing. For each subject, we use all frames from the training data for training, 6 frames with 4 camera viewpoints for novel view synthesis, and all frames with a front camera view for novel pose synthesis. For the DNA-Rendering dataset, we use 416 sequences from Part 1 and 2 for training and 10 sequences for evaluation. For each subject, we use all frames from the training data for training, 48 frames with 4 camera viewpoints for novel view synthesis, and all frames with a front camera view for novel pose synthesis. The foreground masks, camera, and SMPL parameters from these two datasets are used for evaluation purposes.
@@
Comparison Methods. We compare our HumanGif with two categories of state-of-the-art single-view-based animatable human modelling methods, i.e., a generalizable Human NeRF method, SHERF (Hu et al., 2023), and diffusion-based 2D character animation methods, AnimateAnyone (Hu, 2024) and Champ (Zhu et al., 2024). For fair comparisons, we retrain SHERF and Champ by using their official codebase and AnimateAnyone with the open-source implementation from MooreThreads [^1] https://github.com/MooreThreads/Moore-AnimateAnyone.
Implementation Details. The training phase is divided into three stages. In the first stage, we process all frames by resizing and cropping the images to be 512x512 resolution for the RenderPeople dataset and 768x768 resolution for the DNA-Rendering dataset. Then we train a model to learn the mapping from a reference image and a target pose image to a target human image. This stage is trained for 150,000 iterations with a batch size of 4. In the second stage, we incorporate the multi-view attention layer after the spatial-attention and cross-attention in the denoising UNet to learn multi-view consistency. This stage is trained for 20,000 iterations with a batch size of 24 frames (a video length of 6 and 4 views) on each GPU. In the last training stage, we train the temporal layer with the same batch size used in the second stage for 20,000 iterations.
Evaluation Metrics. We evaluate the performance on three metrics, i.e., peak signal-to-noise ratio (PSNR) (Sara et al., 2019), structural similarity index (SSIM) (Wang et al., 2004) and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018). To further evaluate the video fidelity of animatable human videos produced from these methods, we follow (Shao et al., 2024) to report Fréchet Video Distance (FVD) (Unterthiner et al., 2018). As the multi-view RenderPeople released from (Hu et al., 2023) does not contain animatable human videos, we omit the FVD metric in this dataset.
4.2. Quantitative Results
4.3. Qualitative Results
@@
Figure 5. Qualitative results of ablation study on DNA-Rendering dataset.
4.4. Ablation Study
@@
5. Discussion and Conclusion
We propose HumanGif, a single-view conditioned human diffusion model with generative prior. To produce view-consistent and pose-consistent outputs, we incorporate generative prior, a Human NeRF module, multi-view attention, and temporal attention modules into our HumanGif. Furthermore, we introduce an image-level loss during optimization to bridge the gap between latent and image spaces in diffusion models. On two multi-view human datasets, our HumanGif achieves the best perceptual results.
Limitation. 1) There still exist artifacts in our produced results. How to align target images with the input image and target poses remains a future direction to be explored. Meanwhile, we observe that VAE produces distortions for facial areas in some cases. Introducing VAE tailored for human images is also a promising direction to erase artifacts. 2) It is still challenging to ensure a novel view and pose consistency in the long sequence generation. We hope our research could inspire more future research to solve these problems.

### v2

@@ 3.2 Single-view conditioned Human Diffusion Model
SHERF [49] 9.75E-06 26.128 0.934 0.063 7.48E-06 27.435 0.946 0.048
Animate Anyone [47] 2.32E-05 21.022 0.929 0.064 1.35E-05 24.306 0.946 0.041 -
Champ [143] 2.42E-05 21.326 0.930 0.064 1.34E-05 25.381 0.952 0.037 -
@@
SHERF [49] 3.77E-06 27.581 0.948 0.049 3.99E-06 27.471 0.944 0.048 12.01
Animate Anyone [47] 6.20E-06 23.499 0.902 0.056 6.84E-06 23.043 0.907 0.061 16.22
Champ [143] 6.37E-06 23.715 0.869 0.054 6.96E-06 23.253 0.879 0.058 16.89
@@
Human NeRF. The proposed framework aims to generate consistent novel views and poses from a single human image. However, we observe two challenges in the learned human diffusion models: (a) difficulty in capturing fine-grained details (e.g., logos on T-shirts), even when they are present in the reference image, and (b) limited capability to maintain consistency across multiple views and poses. One underlying reason for the challenges is that the reference image is first encoded into a latent representation and then processed by ReferenceNet to extract features, resulting in information loss during this process. To tackle this challenge, inspired by prior research in Human NeRF [89, 49, 19, 87, 125, 85, 118, 109, 59, 61, 114, 70, 33, 23, 139, 52], we integrate a Human NeRF module into human diffusion models. This module learns fine-grained features that are spatially aligned with both the reference and target pose space, enhancing detail preservation and consistency.
As illustrated in Fig. 2, the input to the Human NeRF module is a single human image $\mathbf{I}^{\text{ref}}$ along with its corresponding camera parameters $\bm{P}^{\text{ref}}$ and SMPL pose parameter $\bm{\theta}^{\text{ref}}$ and shape parameter $\bm{\beta}^{\text{ref}}$. The module outputs the rendered human feature image in the target camera view $\bm{P}^{\text{tgt}}$, corresponding to the target SMPL pose $\bm{\theta}^{tgt}$ and shape $\bm{\beta}^{\text{tgt}}$. Specifically, in the target space, we cast rays passing through the camera origin and image pixels, and sample points $\bm{x}^{\text{tgt}}$ along the cast rays. These points $\bm{x}^{\text{tgt}}$ are transformed into the canonical space $\bm{x}^{c}$ using inverse Linear Blend Skinning (LBS). Subsequently, hierarchical 3D-aware features,i.e., point-level features and pixel-aligned features, are queried from their respective feature extraction modules. The queried features are concatenated and passed to the NeRF decoder to predict the density $\sigma$ and feature $\bm{c}$ for each sampled point. The final pixel features are rendered in the target space through volume rendering, integrating the density and feature values of the sampled 3D points along the rays in the target space.
@@
The details of extracting point-level features and pixel-aligned features are described as follows. 1) Point-level Features. We first extract per-point features by projecting the SMPL vertices onto the 2D feature map of the input image, as illustrated in Fig. 2. Next, we apply inverse LBS to transform the posed vertex features into the canonical space. These transformed features are then voxelized into sparse 3D volume tensors and further processed using sparse 3D convolutions [24]. From the encoded sparse 3D volume tensors, we extract point-level features $\bm{f}_{\text{point}}(\bm{x}^{c})$ for each point $\bm{x}^{c}$. With the awareness of 3D Human structure, point-level features capture local texture details in the seen area and infer textural information for unseen areas through sparse convolution. 2) Pixel-aligned Features. Due to limited SMPL mesh and voxel resolution, point-level features suffer from significant information loss, especially in areas visible from the reference image. To compensate for the information loss problem, we additionally extract pixel-aligned features by projecting 3D deformed points $\bm{x}^{c}$ into the input view. As illustrated in Fig. 2, each deformed point $\bm{x}^{c}$ is transformed into the observation space as $\bm{x}^{\text{ref}}=\text{LBS}(\bm{x}^{c};\bm{\theta}^{\text{ref}},\bm{\beta}^{\text{ref}})$ using LBS. It is then projected onto the input view, allowing us to query the pixel-aligned features. Leveraging the complementary strengths of point-level and pixel-aligned features, our Human NeRF module effectively captures fine-grained feature details in regions visible in the reference view while inferring features for regions occluded in the reference view.
@@
Multi-view/Temporal Module. To learn view-consistent and temporal coherent 3D human avatars, we further integrate a view-attention and temporal attention layer after the spatial-attention and cross-attention components within the Res-Trans block of Denosing UNet. Inspired by the efficient temporal attention layer adopted in [47, 35], we utilize the same architecture for our view-attention and temporal attention layer. The view-attention layer performs the attention in the view channel while the temporal attention layer performs the attention in the human pose channel.
3.3 Training Detail
@@
$\mathcal{L}=\mathcal{L}_{\text{diff-latent}}+\lambda_{1}\mathcal{L}_{\text{diff-img}}+\lambda_{2}\mathcal{L}_{\text{NeRF}},$ (4)
@@
To bridge the gap between latent space and image space [83], we formulate $\mathcal{L}_{\text{diff-img}}$ with approximation $\bm{\tilde{z}}_{0}(\bm{\epsilon},\Phi):=\bm{z}_{0}+\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}(\bm{\epsilon}-\bm{\epsilon}_{\Phi}(\bm{z}_{t},t,\bm{c}))$,
$\mathcal{L}_{\text{diff-img}}=w_{t}\mathbb{E}_{\bm{z}_{t},\bm{\epsilon},t,\bm{c}}\Arrowvert\mathcal{D}(\bm{\tilde{z}}_{0}(\bm{\epsilon},\Phi))-\bm{I}^{\text{tgt}}\Arrowvert^{2},$ (5)
@@
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, $\bm{I}^{\text{tgt}}$ is the image at the target view with the target human pose, $\mathcal{D}$ denotes the decoder in a VAE [68, 113] model. Since the approximation holds reliably only for steps $t$ close to zero, we introduce a weight $w_{t}$ that progressively decays as the step value increases. In addition to MSE loss, we apply structural similarity index (SSIM) [117] and Learned Perceptual Image Patch Similarity (LPIPS) [136] as additional image-level loss terms. For features from Human NeRF, we follow [17] to add a term $\mathcal{L}_{\text{NeRF}}$, which regularizes the first three channels by computing MSE, SSIM, and LPIPS with the target image $\bm{I}^{\text{tgt}}$ and a Mask Loss with the target human mask.
4 Experiments
4.1 Experimental Setup
@@
Datasets. We evaluate the performance of our HumanGif on two multi-view human modelling datasets, i.e., RenderPeople [1] (Synthetic Humans) and DNA-Rendering [22] (Real-World Humans). For RenderPeople, we randomly sample 450 subjects as the training set and 30 subjects for testing. For each subject, we use all frames from the training data for training, 6 frames with 4 camera viewpoints for novel view synthesis, and all frames with a front camera view for novel pose synthesis. For the DNA-Rendering dataset, we use 416 sequences from Part 1 and 2 for training and 10 sequences for evaluation. For each subject, we use all frames from the training data for training, 48 frames with 4 camera viewpoints for novel view synthesis, and all frames with a front camera view for novel pose synthesis. The foreground masks, camera, and SMPL parameters from these two datasets are used for evaluation purposes.
Comparison Methods. We compare our HumanGif with two categories of state-of-the-art single-view-based animatable human modelling methods, i.e., a generalizable Human NeRF method, SHERF [49], and diffusion-based 2D character animation methods, AnimateAnyone [47] and Champ [143]. For fair comparisons, we retrain SHERF and Champ by using their official codebase and AnimateAnyone with the open-source implementation from MooreThreads [^1] https://github.com/MooreThreads/Moore-AnimateAnyone.
Implementation Details. The training phase is divided into three stages. In the first stage, we process all frames by resizing and cropping the images to be 512x512 resolution for the RenderPeople dataset and 768x768 resolution for the DNA-Rendering dataset. Then we train a model to learn the mapping from a reference image and a target pose image to a target human image. This stage is trained for 150,000 iterations with a batch size of 4. In the second stage, we incorporate the multi-view attention layer after the spatial-attention and cross-attention in the denoising UNet to learn multi-view consistency. This stage is trained for 10,000 iterations with a batch size of 24 frames (a video length of 6 and 4 views) on each GPU. In the last training stage, we train the temporal layer with the same batch size used in the second stage for 10,000 iterations.
Evaluation Metrics. To quantitatively compare our HumanGif with baseline methods, we evaluate the performance on three metrics, i.e., peak signal-to-noise ratio (PSNR) [102], structural similarity index (SSIM) [117] and Learned Perceptual Image Patch Similarity (LPIPS) [136]. To further evaluate the video fidelity of animatable human videos produced from these methods, we follow [105] to report Fréchet Video Distance (FVD) [112]. As the multi-view RenderPeople released from [49] does not contain animatable human videos, we omit the FVD metric in this dataset.
Figure 3: Qualitative results of novel view synthesis (1st and 3rd row) and novel pose synthesis (2nd and 4th row) produced by SHERF, Animate Anyone, Champ and our HumanGif on RenderPeople and DNA-Rendering datasets.
4.2 Quantitative Results
Table 2: Ablation study on DNA-Rendering. The left side shows different design components that are ablated on.
Generative Prior Human NeRF Module Image-Level Loss Multi-View Attention Motion Attention Novel View Novel Pose
@@ 4.2 Quantitative Results
21.657 0.921 0.069 20.874 0.910 0.076 15.38
$\checkmark$ 23.684 0.860 0.056 23.080 0.870 0.061 17.96
$\checkmark$ $\checkmark$ 24.026 0.923 0.050 24.598 0.923 0.050 11.72
$\checkmark$ $\checkmark$ $\checkmark$ 22.731 0.927 0.050 23.376 0.930 0.049 11.59
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23.647 0.935 0.049 23.504 0.931 0.048 11.24
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23.686 0.935 0.047 24.275 0.935 0.045 9.88
Figure 4: User preference scores on synthesis results.
4.3 Qualitative Results
@@
Figure 5: Qualitative results of ablation study on DNA-Rendering dataset.
4.4 Ablation Study
@@
5 Discussion and Conclusion
To sum up, we propose HumanGif, which reformulates the single-view-based 3D human novel view and pose synthesis as a single-view-conditioned human diffusion process, utilizing generative priors from foundational diffusion models to complement the missing information. To further produce perpetually realistic, view-consistent, and temporally coherent human avatars from a single image, we incorporate a Human NeRF module along with a view-attention, and temporal attention layer into our HumanGif. Furthermore, we introduce an image-level loss during optimization to bridge the gap between latent and image spaces in diffusion models. On two multi-view human datasets, our HumanGif achieves the best perceptual results.
Limitation and Future Work. 1) While our proposed HumanGif improves the synthesis performance, there still exists inconsistency between the generated images and the input image. How to align target images with the input image and target poses remains a future direction to be explored. 2) There is still room to improve the visual quality of generated results. We observe that VAE produces distortions for facial areas in some cases. Introducing VAE tailored for human images is a promising direction to improve the quality. 3) Currently, our HumanGif is trained on existing multi-view human datasets. Scaling our HumanGif to larger human datasets remains a future direction to be explored.